OpenStudy (shamil98):

∫ u dv = uv - ∫ v du ∫ x^6 ln x dx u = x^6 du = 6x^5 dv = ln x v = x ln(x) - x x^6(x ln(x) - x) - ∫ (x ln(x) - x) 6x^5 ?

OpenStudy (anonymous):

you could do it that way. A simplier approach is to let u = ln(x) and dv = x^6

OpenStudy (shamil98):

Ok. so ∫ x^6 ln x dx u = ln x dv = x^6 du = 1/x v = x^7/7 ln(x) * x^7/7 - ∫ (x^7/7)(1/x) ?

OpenStudy (anonymous):

yep

OpenStudy (shamil98):

ln(x) * x^7/7 - ∫ (x^7/7)(1/x) ln(x)* x^7/7 - x^8/56 * ln(x) + C ?

OpenStudy (anonymous):

∫ (x^7/7)(1/x) dx = (1/7)∫x^6 dx = (1/49)x^7

OpenStudy (shamil98):

oh. ln(x) * x^7/7 - ∫ (x^7/7)(1/x) ln(x) * x^7/7 - ∫ x^6/7 ln(x) * x^7/7 - ∫ (x^7/49 my bad didnt simplify that part.

OpenStudy (shamil98):

ln(x) * x^7/7 - x^7/49 + C **

OpenStudy (anonymous):

correct

OpenStudy (shamil98):

ty